OpenStudy (anonymous):

Multiply and simplify: radical 6 (radical 3+5 radical2)

sam (.sam.):

\[3\sqrt{2}+10\sqrt{3}\]

OpenStudy (sasogeek):

.sam ummm u mind showing steps...? i don't understand what you've done

sam (.sam.):

\[\sqrt{6}(\sqrt{3}+5\sqrt{2})\] \[\sqrt{18}+5\sqrt{12}\] \[\sqrt{9}\sqrt{2}+5\sqrt{3}\sqrt{4}\] \[3\sqrt{2}+5(2)\sqrt{3}\] \[3\sqrt{2}+10\sqrt{3}\]